Although I know the answer to the following problem, I need to write it as a function and make a table. I hope you can help me. Here is the problem:
The temperature of a pool is 75 degrees. When the heater is on, the temperature of the pool rises 2 degrees every hour. What will be the temperature of the pool in 3 hours? Write a function and a table to solve this problem.

Answer by stanbon(57361)

The temperature of a pool is 75 degrees. When the heater is on, the temperature of the pool rises 2 degrees every hour. What will be the temperature of the pool in 3 hours? Write a function and a table to solve this problem.
-----------------------
If x = 0, temp = 75
If x = 1, temp = 75+2*1
If x = 2, temp = 75+2*2
If x = 3, temp = 75+2*3 = 75+6 = 81 degrees
-------------------------
If x = x, temp = 75+2x
------------------
EQUATION:
t(x) = 75+2x
=========================
Cheers,
Stan H.